If
(2x+1)(x-5)
----------- =0,then x=
x+3
Note-(2x+1)(x-5)is divided by x+3
: If
(2x+1)(x-5)
----------- =0,then x=
x+3
Note-(2x+1)(x-5)is divided by x+3

Answer by Fombitz(1799)

((2x+1)(x-5))/(x+3)=0
((2x+1)(x-5))=0 Multiply both sides by (x+3).
Solve for the two factors being equal to 0.
1.(2x+1)=0
1.x=-1/2
2.x-5=0
2.x=5